OpenStudy (anonymous):

What is the slope of the line whose function f(-2)=1 satisfies and f(5)=-10

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

I got it... \[\large \color{green}{Slope = \frac{y_2 - y_1}{x_2 - x_1}}\] \[Slope = \frac{-10 - 1}{5-(-2)} = \frac{-11}{7}\]

OpenStudy (anonymous):

Since there is given functions to us so, I can rewrite the formula of Slope as: \[\large \color{red}{Slope = \frac{f(x_2) - f(x_1)}{x_2 - x_1}}\] Here, \(f(x_2) = -10\) \(f(x_1) = 1\) \(x_1 = -2\) \(x_2 = 5\)

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

Careful..